OpenStudy (anonymous):

squareroot X +5 - squareroot x-3=4

OpenStudy (anonymous):

Can I prove in someway that this problem dosnt have a solution?

OpenStudy (anonymous):

\[\sqrt{x+5} - \sqrt{x-3} = 4\]